ID |
Date |
Author |
Type |
Category |
Subject |
3372
|
Fri Aug 6 10:44:55 2010 |
nancy, koji | Update | IOO | Mode Cleaner WFS |
Quote: |
Upon Nancy's request, I checked the status of the suspensions.
I found that the power strip of the 1Y4 rack was turned off.
Since it has a over current breaker, I don't know whether it happened by someone or over current.
Anyway, I restarted the sus computers, and now the suspensions are damping as usual.
The MC has been aligned, the auto locker is also working.
Incidentally, I found that the WFS servos are not working. Actually since the last night
It repeated losing lock and unlock.
Probably some values of the matrix or the gain is wrong.
I left the WFS as it is because Nancy will put new values this afternoon.
I will ask her to confirm that the old values work at the end of her work.
|
Yesterday , I put in the Output Matrix, and changed the gain sliders for the 4 WFS loops.
It worked and was keeping the lock for the MC.
I then tested whether the MC1 and 3 were following any change in MC2 alignment. It was indeed workinng,
Next we stepped to putting in the gains for the MC2 oplev servo.
the signs are decided on the basis of convergence, and the magnitude is kept very low, to have a very slow control for MC2.
This complete 6 * 6 model does work, and was able to keep the transmission held.
I also tried poking each mirror in pitcg and yaw, and the cavity comes back to high resonance after some time.
This time is indeed large if the poking is made for MC2, and the transmission comes back to normal after big oscillations.
I tried to measure the Open loop TFs for all these loops yesterday, but somehow could not find a correct excitation.
I will do it today.
Plan ahead :
1. Center the spot on MC2 and the QPD
2. Optimize the gains by looking at response to noise.
3. Measure Power Spectrum Density of each error signal.
|
3371
|
Fri Aug 6 08:09:15 2010 |
steve | Update | PSL | PSL preparation work |
Quote: |
PSL preparation work
Aug 5th 10am-4pm?: (Kiwamu, Alberto, Koji)
Removing the unused cables around the PSL table and the control room
Aug 6th 10am-4pm?: (Kiwamu (ex. noon-2pm), Alberto, Koji, Jenne ('till noon) )
- Labeling the cables to be disconnected / making the records ==> All
- Removals
- the big nuts at the side of the PSL table ==> Steve
- the cable ties on the PSL frame ==> Easy
- Innolight 2W ==> Kiwamu
- the green pickoff optics at the edge, if necessary ==> Kiwamu talking with Steve
- Optics on the shelf ==> Jenne / Koji
- Oscilloscopes on the shelf ==> Jenne / Koji
- CCD camera connections (optional, as far as not critical for the operation)
- Put poles on the table (for the plastic sheet) ==> Alberto / Koji
Aug 9th 9am-5pm: (Steve, Jenne, Alberto, Koji)
- Disconnecting the cables ==> All
- Shutting down the PSL ==> Steve/Koji
- Draining the cooling water ==> Steve
- Removals
- The accelerometers ==> Jenne
- the PSL chamber ==> Steve
- Periscopes ==> Alberto
- Sealing of the table with the plastic sheets
- The chiller is planned to go to MIT
|
Monday, August 9
We should move the reference cavity too. Will this cavity be pumped while relocated?
Check and insure that attached and cut-free cables of PSL have enough room to tolerate the raising of the enclosure by 6"
I had second thoughts about the power line to the OMC. Koji was right, we should disconnect them from the power supplies.
The PSL enclosure doors on the north side will have to be removed some times to move exiting and entering ports.
|
3370
|
Thu Aug 5 22:36:11 2010 |
Koji | Update | PSL | PSL preparation work | PSL preparation work
Aug 5th 10am-4pm?: (Kiwamu, Alberto, Koji)
Removing the unused cables around the PSL table and the control room
Aug 6th 10am-4pm?: (Kiwamu (ex. noon-2pm), Alberto, Koji, Jenne ('till noon) )
- Labeling the cables to be disconnected / making the records ==> All
- Removals
- the big nuts at the side of the PSL table ==> Steve
- the cable ties on the PSL frame ==> Easy
- Innolight 2W ==> Kiwamu
- the green pickoff optics at the edge, if necessary ==> Kiwamu talking with Steve
- Optics on the shelf ==> Jenne / Koji
- Oscilloscopes on the shelf ==> Jenne / Koji
- CCD camera connections (optional, as far as not critical for the operation)
- Put poles on the table (for the plastic sheet) ==> Alberto / Koji
Aug 9th 9am-5pm: (Steve, Jenne, Alberto, Koji)
- Disconnecting the cables ==> All
- Shutting down the PSL ==> Steve/Koji
- Draining the cooling water ==> Steve
- Removals
- The accelerometers ==> Jenne
- the PSL chamber ==> Steve
- Periscopes ==> Alberto
- Sealing of the table with the plastic sheets
- The chiller is planned to go to MIT
|
3369
|
Thu Aug 5 17:59:23 2010 |
Koji | Update | PSL | Cable removal from the control room | [Alberto, Kiwamu, and Koji]
We removed the BNC cables from the control room.
The work was as hard as the one I had when I swept a 300m tunnel with a vacuum...
If we could remove the video cables, that would be a real epoch.
We found that the cabling behind the AP table is still quite ugly....grurrrh |
Attachment 1: IMG_2684.jpg
|
|
3368
|
Thu Aug 5 16:18:37 2010 |
Alastair | Update | General | clocks still not locked fully - hut removed | The rubidium clocks are still not quite locked together, though it is clear that the beat frequency has dropped a lot since yesterday.
I checked on the clocks and the 1pps sync light is on. The clocks are really hot again though despite the gap left at the bottom of the igloo. The side of the clocks were hot enough to not be touchable.
I made the executive decision that I would remove the hut just now. We can let the clocks lock together and then put the hut back on just before measuring. This way the hut will isolate from temperature fluctuations during the measurement but it won't be running at the hotter equilibrium temperature. I hope that the temperature won't change too much during the measurement if we put the hut back on.
RA: Attachment deleted because it was in Postscript format. Also not allowed here are the Stone Tablet and Cave Painting formats. |
3367
|
Thu Aug 5 13:05:53 2010 |
Koji | Update | IOO | Mode Cleaner WFS | Upon Nancy's request, I checked the status of the suspensions.
I found that the power strip of the 1Y4 rack was turned off.
Since it has a over current breaker, I don't know whether it happened by someone or over current.
Anyway, I restarted the sus computers, and now the suspensions are damping as usual.
The MC has been aligned, the auto locker is also working.
Incidentally, I found that the WFS servos are not working. Actually since the last night
It repeated losing lock and unlock.
Probably some values of the matrix or the gain is wrong.
I left the WFS as it is because Nancy will put new values this afternoon.
I will ask her to confirm that the old values work at the end of her work. |
3366
|
Thu Aug 5 11:48:52 2010 |
nancy | Update | IOO | Mode Cleaner WFS |
Quote: |
I calculated the MC1&3 Vs WFS1&2 Output Matrix today from the above measurements with koji's help.
the matrix can be generated from the m file at /cvs/cds/caltech/users/nancy/Align_Matrix/matrix.m
these values were put in, and the direction of control is sort of confirmed. I tried twiddling with the gains in the loop to get a 4*4 stable control, but could not succeed.
the mode cleaner is back locked now, and WFS matrix as well as gains are reverted to the old values. (1.30 am)
The output Matrices are
Pitch
0.724
|
0.197
|
0
|
0
|
-1.448
|
-0.758
|
Yaw
0.919
|
0.139
|
0
|
0
|
-0.0106
|
-0.1245
|
|
I realised today morning that there was a flaw in my calculations for the yaw matrix.
Correcting the values, and also making teh tables more readable.
I will test these values once our computers are back to working condition.
PITCH
|
WFS1
|
WFS2
|
MC1
|
0.724
|
0.1964
|
MC2
|
0
|
0
|
MC3
|
-1.4436
|
-0.756
|
YAW
|
WFS1
|
WFS2
|
MC1
|
0.0710
|
0.1074
|
MC2
|
0
|
0
|
MC3
|
0.0082
|
-0.962
|
|
3365
|
Thu Aug 5 01:29:39 2010 |
nancy | Update | IOO | Mode Cleaner WFS |
I calculated the MC1&3 Vs WFS1&2 Output Matrix today from the above measurements with koji's help.
the matrix can be generated from the m file at /cvs/cds/caltech/users/nancy/Align_Matrix/matrix.m
these values were put in, and the direction of control is sort of confirmed. I tried twiddling with the gains in the loop to get a 4*4 stable control, but could not succeed.
the mode cleaner is back locked now, and WFS matrix as well as gains are reverted to the old values. (1.30 am)
The output Matrices are
Pitch
0.724
|
0.197
|
0
|
0
|
-1.448
|
-0.758
|
Yaw
0.919
|
0.139
|
0
|
0
|
-0.0106
|
-0.1245
|
|
3364
|
Thu Aug 5 00:17:41 2010 |
Koji | Update | PSL | PSL preparation work | We start the work on the cables at around the PSL table.
Aug 5th 10am-4pm?: (Kiwamu, Alberto, Koji)
- Removal of the unused cables around the PSL table and the control room
- Removal of the cable ties on the PSL frame
- Removal of the big nuts at the side of the PSL table
Aug 6th 10am-4pm?: (Kiwamu, Alberto, Koji, Jenne (~noon) )
- Labeling of the cables
- Planning of the disconnection
Aug 9th 9am-5pm: (Steve, Jenne, Alberto, Koji)
- Shutting down of the PSL
- Disconnection of the cables
- Draining of the cooling water
- Removal of the accelerometers
- Removal of the PSL chamber
- Sealing of the table with the plastic sheets |
3363
|
Wed Aug 4 20:58:22 2010 |
Gopal | Update | WIKI-40M Update | 7.28.10 - 8.4.10 Weekly Update | Summary of this week's activities:
7/28: Finished Y-Translational 4-Stack Analysis
"Tapered Cantilever" COMSOL tutorial
Tried (and failed) isolating gravity from oscillation
7/29: Developed tilt/rotation load combinations for torsional inputs and showed these to work in the model
Tried using Normal Vector mode on top plate to obtain output tilts; worked for the rectangular bar, but not for the full stack 
Talked to Jan about a 1st-order alternative to gravity - requires Weak Form (only found in COMSOL 3.5 right now)
Began Z-Translational 4-Stack Analysis -- Ran Overnight
7/30: Progress Report 1st Draft
Completed Z-Translational 4-Stack Analysis
8/1: Progress Report 2nd Draft
8/2: Progress Report 3rd Draft
Submitted Progress Report
8/3: Finalized Eigenfrequency Analysis for MC1/MC3 Stack 
24 Physical Eigenmodes plotted and recorded, as expected
Should be good enough for the final report --> focus on transfer function analysis for the remainder of the SURF
8/4: Prescribed Displacement Tests on Simple Rectangular Block --> shown to better produce displacement-displacement transfer functions
X-to-X Transfer Function seems much better when plotted
Should now be able to do the Displacement portion of Transfer Function Analysis on MC1/MC3 for Translational Modes
(I apologize that this update is a little late) |
3362
|
Wed Aug 4 20:15:07 2010 |
Jenne | Update | PEM | Yay! Guralps work again! | After much hassle, the Guralp cable from the ADC Out of the breakout box to the ADC is fixed, and everything is plugged in and working again. The seismometers are back in their regular positions at the ends of the MC, ready for some excellent seis/MC combo data.
I solidified the change of putting the Gur2Z channel into a different BNC input on the ADC. The C1ADCU_PEM.ini file has been changed so that what used to be the Ranger's channel is now recognized as Gur2Z.
Also, I changed the same .ini file to reflect Koji's move of ACC_MC1_Z to the old AUDIO_MIC2 channel, so now all 6 Accelerometer channels have the same calibrations again.
Another big change is the change from old-left-handed convention to new-right-handed convention. The seismometers are aligned the same way they always have been (with the North-South markers aligned with the MC), but now the North-South output is plugged into the BNC on the ADC that is associated with Gur*_X, and the East-West output is plugged into the ADC channel associated with Gur*_Y. This is true for both Guralp Seismometers.
So, now we have:
Gur1_X = Gur1_NS = ADC#10
Gur1_Y = Gur1_EW = ADC#11
Gur1_Z = Gur1_Vert = ADC#12
Gur2_X = Gur2_NS = ADC#2
Gur2_Y = Gur2_EW = ADC#3
Gur2_Z = Gur2_Vert = ADC#24
SEIS_Ranger_Y = no longer in the .ini file |
3361
|
Wed Aug 4 19:50:58 2010 |
rana | Configuration | Electronics | Rubidium clocks too hot: hut removed | Alastair found that the foam hut that he and Jan put on top of the Rb clocks to temperature stabilize them was too good of an insulator. The Rb boxes had gotten very hot and became internally unlocked as seen on the front panel.
After we let them cool down with the box off, I turned them back on. After several minutes the 'Locked' light came back on. Some minutes after that the '1PPS Sync' light also came on, indicating that the two had become somewhat synchronized. It really means that the frequencies are kind of close: I think its roughly that f1-f2 < 2 mHz.
I put the yellow box back on and have left it with a small gap on the bottom so that the hot air can get out. Hopefully, this will protect the clocks from the wind, but not cause them to overheat.
The signal going to the DAQ right now is DC-coupled, with a gain of 1. The peak-peak beat signal in this situation is 6300 counts.
My guess is that the clocks will by synchronized by tomorrow afternoon so that we can get the measurement done. Please don't disturb the clocks or the yellow box around them. Try to minimize any activity around that area. |
Attachment 1: beat.png
|
|
3360
|
Wed Aug 4 16:52:59 2010 |
Razib, Aidan | Update | Phase Camera | Sideband power measurement (updated) | Aidan and I made some attempt to measure the power of the sidebands so that we can calculate our expected signal strength.
Our setup looks like the following:

As light from the laser is split into two at BS1, the transmitted beam has higher power as our BS1 is only coated for 1064nm. We get two reflected beams from BS1, one reflected of the front surface and the other from the back surface. We took the stronger back reflected beam to the EOM driven at 40 MHz (also at 25 MHz at a later time). The AOM produced a reference beam with 40 .000 005 MHz offset which we recombined with the sidebands obtained from the EOM. The beat produced is sent off to PDA 10CF connected to 4395A spectrum analyzer.
The plots for 40MHz sidebands and 25 MHz sidebands looks like this:

From the above spectra, at 40 MHz sideband regime:
Power of the carrier @ 40 MHz = -39.72 dBm
Power of the sideband @ 80 MHz = -60.39 dBm

At 25 MHz sideband regime,
Power of the carrier @ 40 MHz = -40.22 dBm
Power of the upper sideband @ 65 MHz = -61.72 dBm
Power of the lower sideband @ 15 MHz = -60.99 dBm
Power Measurement:
We made some necessary power measurement using a PD connected to a voltmeter after the EOM and the AOM when the EOM is driven at 40 MHz:
___________________________________________________________
Dark : 0.025 V
AOM on: 4.10 V (EOM blocked)
EOM : 2.425 V (AOM blocked)
___________________________________________________________
From the earlier calculation (ref: Elog entry July 28) the power that we expect to see at the PD is,
P= A_c ^2 + A_r^2 + A_(-sb)^2+ A_sb ^2 +2* A_r* A_sb * cos ( w_(r,sb) t ) , where A_c= carrier; A_r= reference beam; A_sb=Upper sideband; A_(-sb)= Lower sideband, w_(r,sb) = w_r - w_sb
P = A_c ^2 + A_r^2 + A_(-sb)^2+ A_sb ^2 +2* A_r* A_sb , letting cos (w_(r,sb) go to 1) is order to approximate the maximum signal
So the signal that we expect to see relative to the DC ( i.e A_c ^2 + A_r^2 + A_(-sb)^2+ A_sb ^2, the first four terms of the power equation) is,
Sig = 2* A_r* A_sb / { A_c ^2 + A_r^2 + A_(-sb)^2+ A_sb ^2 },
Since the modulation index is small, the power in the sideband is very small compared to carrier and the reference beam. So we can ignore the sideband power for the signal expression.
So,
Sig = 2* A_r* A_sb / ( A_c ^2 + A_r^2 )
So if we want to maximize this signal w.r.t the reference then,
d (sig)/ d(A_r) = 2 { ( A_c ^2 - A_r^2) *A_sb } / {( A_c^2 + A_r^2)} ^2
Thus, the signal is maximized when,
A_r^2 = A_c^2
We adjusted the AOM to be driven at + 7.7 dBM so that the new power at the AOM matched the EOM power, which is 2.397 in the voltmeter.
So the power at both the AOM and the EOM are:
P_AOM = ( V_AOM - V_dark) / (PD responsitivity * Transimpedance gain)
= ( 2.397 - 0.025 ) / ( 0.45 * 1.5 x 10 ^5 )
= 3.51 x 10 ^ - 5 W
P_EOM = (V_EOM - V _dark) / (PD responsitivity * Transimpedance gain)
= ( 2. 425 - .0.025) / ( 0.45 * 1.5 x 10 ^5 )
= 3.55 x 10^ - 5 W
From the spectra of the 40 MHz sideband above, the ratio of the carrier and the sideband amplitude is: A_c / A_sb = 10.8 .
P_EOM = A_c ^2 + 2 A_sb ^2
Therefore, A_sb = sqrt ( P_EOM / 118.64) = 5.47 x 10^ - 4 V/m
Thus, A_c = 5.908 x 10^ -3 V/m
and A_r = sqrt ( P_AOM) = 5.92 x 10 -3 V/m.
This measurement can be used to calculate the signal to contrast ratio (SCR) that we expect to see:
SCR = 2 A_r * A_sb / ( A_c^2 + A_r^2 ) = 0.09
Our next step is to measure the actual signal to constrast ratio as seen by the camera. Details of that will be posted soon. |
3358
|
Wed Aug 4 12:49:42 2010 |
nancy | Update | IOO | Mode Cleaner WFS |
Quote: |
Yesterday, I started twiddling with the Mode Cleaner at about 2 pm.
So the seismic data should be all good before that.
I was using it till about 3.30 am, and then left for the night with locking it and swithcing on back the WFS control
Today morning, I have started twiddling with it again, at about 10.30 am.
About my work with the mode cleaner :
I am primarily exciting the mirrors in pitch and yaw, and trying to measure the response of the WFS and the MC2 OPLEV wrt the excitation.
This thus involves switching off the WFS control while measurement.
After two more of those measurements today, I will get to finding new values for the Output Matrix of the WFS for controlling MC1 & 3, and also, try giving in control to MC2 alignment using OPLEV signals.
|
TFs after the measurement -
In the order - MC1 , MC2 , MC3 -pitch and yaw.
These plots let us know about how do the wavefront sensor signals actually respond to the mis-alignments in the mirrors.
For legibility, legend has been includded in only one plot in each pdf., its typically the same for all 3 plots.
the actual xml files for this measurement are in the directory /cvs/cds/caltech/users/nancy/Align_Matrix/highpower/spot_center
It was made sure before each measurement that the MC is best aligned, the WFS are turned off, and the spots on all 3 QPDs are centered.
|
Attachment 1: pit.pdf
|
|
Attachment 2: yaw.pdf
|
|
Attachment 3: pit.pdf
|
|
Attachment 4: yaw.pdf
|
|
Attachment 5: pit.pdf
|
|
Attachment 6: yaw.pdf
|
|
3357
|
Wed Aug 4 11:10:28 2010 |
nancy | Update | IOO | Mode Cleaner WFS | Yesterday, I started twiddling with the Mode Cleaner at about 2 pm.
So the seismic data should be all good before that.
I was using it till about 3.30 am, and then left for the night with locking it and swithcing on back the WFS control
Today morning, I have started twiddling with it again, at about 10.30 am.
About my work with the mode cleaner :
I am primarily exciting the mirrors in pitch and yaw, and trying to measure the response of the WFS and the MC2 OPLEV wrt the excitation.
This thus involves switching off the WFS control while measurement.
After two more of those measurements today, I will get to finding new values for the Output Matrix of the WFS for controlling MC1 & 3, and also, try giving in control to MC2 alignment using OPLEV signals.
|
3356
|
Wed Aug 4 03:01:56 2010 |
Jenne | Update | PEM | Catastrophic Cable Failures |
Quote: |
After some cable swapping this morning, I have determined which cable is bad. It's the Gur1 cable between the seismometer and the breakout box. This is a milspec -> 37pin d-sub cable. I'll pull out the cable and have a look at it after lunch.
|
So, I was wrong about which cable it was, probably in my rush to get some lunch. The actual culprit was the octopus cable that Bob made waaay back in the day (~2 years ago) to go from the "ADC out" of the breakout box (37pin Male Dsub) to 9 BNCs. As it turns out, the Gur1Z channel of that cable was broken on both ends!!!
On one end, we have the 37 pin Dsub. The cable used was so thick (way too thick for this application) that it made a super rigid connection between the wires and the connector, and any bending of the cable stressed this connection, despite the strain-relief of the backshell. The Gur1Z connection snapped off when I was gently wiggling the connections to check them out. Also, since the wires were all so thick, they didn't really fit into the hole in the backshell, so 2 or 3 of them were squished.....straight through the insulation so that several channels were shorted together / potentially shorted to ground. This may explain some of the nasty behavior that Rana and I had seen (although I might have forgotten to elog? My bad.) that even with the inputs of the breakout box all terminated, there was high coherence between the channels. Terminated inputs should give random noise, so this was fishy.
On the other end of the cable we have the 9 BNCs. I had finished redoing the 37pin end of the cable, and was 'beeping' it to check it out, when to my dismay I found that the Gur1Z channels (the inside and the outer shield of the BNC connector) were shorted together. I removed these 2 wires from the Dsub connector to confirm that the BNC was at fault. Koji looked at the BNC with me after I chopped it off of the cable. Bad news strikes again. To get the wires to fit in the inner pin of the BNC connector, the cable-maker had cut off several strands of the wire to make it skinnier. It appears that over the years these cut-off strands wiggled their way to touching the outer shield. This appears to be a danger for all of the BNCs on this cable: a little bit of torque (which one might expect during plugging and unplugging a BNC) and the 2 sides of the differential measurement will be shorted together.
I then decided to start afresh and make my own cable. I found some AWG26 8-twisted-pair cable laying around underneath the Yarm (since this was all I could find, I was just going to do the Gur1 and Gur2 channels, and leave out the Gur3's). The 37 Dsub side was easy, but I seem not able to connect such skinny wire to the BNC connectors in a robust way. Since this bad cable has so far cost me ~2.5 16-hour days of grief, I don't want my new version of it to also be bad. At this point, I await the advice of one wiser than I. I think BNC connectors are designed for something a little closer to ~20AWG, but I could be wrong. Also, they are obviously optimized for coax cable. So what I have now is never going to be great. Maybe tomorrow I can go to the Electronics Shop / Store and buy BNC connectors that are meant to be soldered-to. That would be awesome.
Since I currently have no functional cable to go from Breakout box to ADC, the Guralps are unplugged for tonight.
Conclusions for the day / evening: Frank, Alastair and Jenna are mostly absolved of blame, although the traveling to Bridge and opening and closing the box (which usually involves more plugging and unplugging of cables) probably didn't help this cable out too much. Also, Bob definitely owes me a Sugar Napoleon or something.
In other news, since the Gur2Z ADC channel is totally wacked, I have taken over (but not renamed) the Ranger channel for Gur2Z for now. Jan still has the Ranger hostage over in Bridge, so this is okay for now. |
3355
|
Wed Aug 4 01:44:56 2010 |
Koji | Configuration | PEM | Accelerometer located on and below the PSL table | [Jenne and Koji]
In order to characterize the seismic vibration of the PSL table, we put the accelerometers on and below the PSL table.
On the PSL: MC2 Accelerometers (X, Y, Z) - being connected to CH1-3 of the preamps
East side of the PSL table. X, Y, and Z is directed to North, East, and Up.
On the ground: MC1 Accelerometers (X, Y, Z) - being connected to CH4-6 of the preamps
Beneath the West side of the PSL table. X, Y, and Z is co-aligned to the MC2 ACC.
I found that the C1:PEM-ACC_MC1_Z has large noise in the low freq (~1Hz) region. I tracked down the noise source
and found the noise is still present in the down stream even when the CH17 (C1:PEM-ACC_MC1_Z ) of the ADC IF BNC
(@1Y7) was terminated.
I consulted with Jenne and decided to connect this channel to CH14, which is vacant and has name
C1:PEM-AUDIO_MIC2 (16k).
(for the details of the channel configurations, see /cvs/cds/caltech/chans/daq/C1ADCU_PEM.ini )
I first tried magnetometer channels to steal, but they didn't seem reacting (and the connected to the wrong channels).
I am feeling that we should once entirely check the I/F box.
Note that there looks the difference of the gain by x10 between C1:PEM-ACC_MC* channels and C1:PEM-AUDIO_MIC2 .
[From Jenne: The gain difference is because the C1:PEM-ACC_MC* channels have gain=10 in the .ini file, while C1:PEM-AUDIO_MIC2 uses the default gain=1. ] |
Attachment 1: IMG_2675.jpg
|
|
Attachment 2: IMG_2678.jpg
|
|
Attachment 3: IMG_2682.jpg
|
|
3354
|
Tue Aug 3 11:50:16 2010 |
Jenne | Update | PEM | Seismometer Problem Tracked down | After some cable swapping this morning, I have determined which cable is bad. It's the Gur1 cable between the seismometer and the breakout box. This is a milspec -> 37pin d-sub cable. I'll pull out the cable and have a look at it after lunch. |
3353
|
Tue Aug 3 11:17:10 2010 |
kiwamu | Update | CDS | Diagrams for Cables needed for CDS test | Current Wiring Setup for the Suspension Controls

New Wiring Plan for the Suspension Controls with the New CDS

Missing Stuff for the CDS test
Ideally we can reuse the existing cables, but some of them may not be long enough for the new wiring.
The diagram below shows extremely non-ideal case.

Some more information will be summarized on the wiki later.
|
3352
|
Tue Aug 3 03:15:06 2010 |
nancy | Update | IOO | MC back to locked mode | I turned the WFS gain to 0.02 back, and the MC is locked, the data for the seismic motion might be meaningful nowforth. |
3351
|
Tue Aug 3 01:51:01 2010 |
Jenne | Update | PEM | Seismometer Update: Still not good, but perhaps getting better |
Quote: |
Today's seismometer diagnosis activities are still underway, this is just an update (since I did some reboots):
Problem 1: X and Z channels on both seismometers were flipped. I unplugged an X cable (East/West on the cable labels) and the Z channel (vert) would go to floating ADC zero. Rana suggested that the ADCs sometimes have random channel hopping, and that a reboot of the c0dcu1 computer which handles the PEM ADCU should fix this problem. I keyed the c0dcu1 / c0daqawg crate, those computers came back just fine, and the channels were no longer flipped. This is a good thing. Although now it's actually the Z channels that were / are bad on both seismometers, not the X's.
While rebooting those computers, c1iovme, c1sosvme, c1susvme1 and c1susvme2 crashed. I rebooted them, although for the first few power cycles c1susvme1&2 couldn't mount /cvs/cds/caltech. Eventually they did, and life is good again. Except that the seismometers are still funny.
|
Some more progress, but still not complete:
Jan and I looked at all of the Gur channels on a 'scope (battery, so as not to be grounded), and 5 of the 6 looked good. We were looking at the BNCs just as they go into the ADC. The one which still looks bad is Gur1Z. The 'scope just doesn't see any signal on that channel.
In addition, the ADC's BNC input #4 (which normally has Gur2Z) looks totally shot. When it's floating, the signal on dataviewer definitely doesn't look floating. I'm probably going to have to move over to another channel, and just give that one up (this ADC already has several channels which have been declared bad, so maybe it's not a surprise that this can happen?)
Since one of the Gur signals looks bad (Gur1Z) and one of the ADC channels looks bad (usual Gur2Z), I switched the Z channels on the ADC board, so the channel being saved as Gur1Z is in fact Gur2Z. This is valid as of ~1:15am until further elog notice.
During my investigations into why Gur1Z is funny, I also looked at the signal on the BNC octopus cable coming straight from the output of the Guralp Breakout Box (this is the cable which goes from "ADC Out" on the back of the box which is a 37 pin D-sub to 9 differential BNCs), and sometimes I saw zero on the 'scope, but sometimes there was a signal which would coincide with jumping tests. Whenever there was a signal however, it was always a way lower amplitude (at least by a factor of 10?) than the other channels.
All of this craziness led to me pulling the Guralp box to investigate.
Upon opening the box, I recalled that the channels go in order: Vert, NS, EW. The Gur1Z channel is that first vert channel, and it's the one which always had a blue input capacitor rather than a surface mount one. Being suspicious of Frank and Alastair, since they seemed unhappy with my capacitor choices, I wondered if they had wiggled the blue cap, and tore something loose. Just in case, and to make things seem more uniform, I replaced the blue cap with a surface mount 1uF cap. (Actually its 0.909uF, replacing the 0.905uF blue cap, according to the black DMM that measures capacitance.) While I was in there, since it had been a problem in the past (elog 2811), I relflowed the solder on some of the resistors, especially near the output op amp.
Anyhow, none of that may have been necessary. All 6 of the Gur channels were examined on a 'scope, using clip doodles to measure the various Test Points on the circuit. I looked at all of the TPs in Gur1Z, and I didn't find that any particular stage was any noisier than the others. Also, all 6 of the Gur channels seemed totally fine in terms of sending a good signal to the output of the box, including Gur1Z which is currently under investigation. All of the channels passed the "output looks ~20x the input" test, and for approximately equal thumping on the ground all 6 channels seemed to have similar amplitude outputs. The Z channels on both channels one and channels two were a little bigger than the X's or Y's, but the 2 Z channels were about the same. This test was done using Guralp2 and the Gur2 cable on both channels 1 and 2, and then checked with Guralp 1, using the Gur2 cable on channels 1. The Vert1 channel always seemed good.
I now am suspicious of one or more of the cables: either the Gur1 cable from the seismometer to the box, or the Vert1 channel of the octopus cable. I'm satisfied that the BNC cables running through the cable tray are okay (although it might not hurt to check that they all successfully send a sine wave...) I opened up the backshell of the Gur1 cable, on the end that connects to the breakout box. Nothing seemed amiss. I still need to Beep the cable to check its connections, and look at the octopus cable.
Recap / Current Status: Breakout Box is reinstalled, both seismometers hooked up. The Z channels on the seismometers are swapped at the ADC input. The dataviewer channels Gur1_X, Gur1_Y, Gur1_Z (which is actually Z of Gur2 seismometer) and Gur2_X, Gur2_Y are all good. Nancy is going to leave the MC in a happy place, and note the time when she's done. Tomorrow I'll check out the cables for the Gur1Z seismometer channel. |
3350
|
Mon Aug 2 21:52:57 2010 |
Koji | Update | IOO | MC is running at the full power | [Nancy and Koji]
We restored the full power operation of the MC.
Restoration of the suspensions
- Found the suspension watch dogs are left turned off.
- Found c1susvme1/2 were not running.
- Launched the realtime processes on c1susvme1/2 and c1iscey
- Restored the watch dogs. The suspensions looked fine.
Preparations for the high power
- Put an ND2.0 before the MCT CCD. Confirmed the ND reflection is damped.
MCT QPD is not necessary to be touched.
The high power operation of the MC / post lock adjustment
- Locked the MC under the autolocker being disabled.
- Adjusted the aperture on the MC2 face camera
- Adjusted the spot positions on the WFS QPDs
- Reverted the scripts to the high power ones
(mcup / mcdown / autolockMCmain40m)
- Logged in to op340m and restarted autolockMCmain40m
The autolocker seems working correctly. |
3349
|
Mon Aug 2 18:02:46 2010 |
Jenne | Update | PEM | Seismometer Update: Still not good, but perhpas getting better | Today's seismometer diagnosis activities are still underway, this is just an update (since I did some reboots):
Problem 1: X and Z channels on both seismometers were flipped. I unplugged an X cable (East/West on the cable labels) and the Z channel (vert) would go to floating ADC zero. Rana suggested that the ADCs sometimes have random channel hopping, and that a reboot of the c0dcu1 computer which handles the PEM ADCU should fix this problem. I keyed the c0dcu1 / c0daqawg crate, those computers came back just fine, and the channels were no longer flipped. This is a good thing. Although now it's actually the Z channels that were / are bad on both seismometers, not the X's.
While rebooting those computers, c1iovme, c1sosvme, c1susvme1 and c1susvme2 crashed. I rebooted them, although for the first few power cycles c1susvme1&2 couldn't mount /cvs/cds/caltech. Eventually they did, and life is good again. Except that the seismometers are still funny. |
3348
|
Mon Aug 2 17:12:28 2010 |
Koji | Update | General | Lab schedule for the week of Aug. 2 | Aug
2 Mon - 5 Thu WFS work (Nancy)
2 Mon - 4 Wed
Jenne: Seismometer fix / Seismic measurements on the PSL table
TT characterization (with Koji)
Preparations ETM suspensions (optional: may be in later weeks)
Kiwamu: CDS test for SUS (may be involving Koji)
Alberto: RF system prep.
All: For 5th and 6th: PSL cabling works Koji
5 Thu PSL Table prep
6 Fri PSL Table prep / Likely to shut down the PSL
9 Mon PSL Table prep / shutting down of the PSL (optional)
10 Tue PSL box Frame lifting
12 Thu PSL table tapping
16 Mon - 17 Tue concrete pouring preparation
19 Thu - 23 Fri Tripod placement
24 Tue - 26 Thu concrete pouring |
Attachment 1: PSL_work_schedule.pdf
|
|
3347
|
Mon Aug 2 10:26:15 2010 |
steve | Configuration | VAC | very slow pumpdown completed |
Quote: |
Bob and Steve closed BS chamber with the help of the manual Genie lift and the pump down started. The PSL shutter was closed and manual block was placed in the beam path. High voltage power supplies were checked to be off.
Pumping speed ~ 1 Torr/min was achieved at 1/8 of a turn opened roughing valve RV1
|
3 days pump down #69 completed. Thanks to Koji and Kiwamu who did more roughing on Saturday and Sunday. |
Attachment 1: veryslowpd.jpg
|
|
3346
|
Sun Aug 1 21:40:27 2010 |
rana | Summary | PSL | FSS: SLOWDC response | I bet you thought that the NPRO slow actuator response could be well represented by a pole at ~0.1 Hz? Well, that's just what they want you to believe.
I attach the response measured in FSS-FAST (with no feedback to the SLOW actuator) when the SLOW is given a step. As you may remember from
kindergarten, the response of a single pole low pass should just be an exponential. Clearly, there's more here than 1 pole.
I also inserted a factor of 0.01 in the FSSSlowServo code so I could make the gain sliders have reasonable values (they used to all be ~1e-3). The SVN and the MEDM snapshot are updated. |
Attachment 1: Untitled.png
|
|
3345
|
Sun Aug 1 21:04:45 2010 |
rana | Summary | Computer Scripts / Programs | MC Autolocker fixed | Someone had left a typo in the MC autolocker script recently while trying to set the lock threshold to 0.09. As a result, the autlocker wouldn't run.
I repaired it, made a few readability improvements, and checked in the new version to the SVN. If you make script changes, check them in. If you think its too minor of a change for a SVN checkin, don't do it at all. |
3344
|
Sun Aug 1 18:50:35 2010 |
Koji | Update | VAC | P1 = 0.48 torr | I resumed the pumping today. Now the pressure is 0.48 torr. The pump was stopped.
Quote: |
I resumed the pumping from 19:00.
Now the valve RV1 is full open. But the pumping is really slow as we are using only one RP.
After 3hrs of pumping, P1 reached 1.2mmtorr but still we need 2hrs of pumping...
I stopped pumping at 22:30.
|
|
3343
|
Sat Jul 31 22:35:01 2010 |
Koji | Update | VAC | Vac-P1 still 1.2mtorr | I resumed the pumping from 19:00.
Now the valve RV1 is full open. But the pumping is really slow as we are using only one RP.
After 3hrs of pumping, P1 reached 1.2mmtorr but still we need 2hrs of pumping...
I stopped pumping at 22:30. |
3342
|
Sat Jul 31 17:37:36 2010 |
josephb | Update | CDS | Cables needed for CDS test | Last Thursday, Kiwamu and I went through the cabling necessary for a full damping test of the vertex optics controled by the sus subsytem, i.e. BS, ITMX, ITMY, PRM, SRM. The sus IO chassis is sitting in the middle of the 1X4 rack. The c1sus computer is the top 1U computer in that rack.
ADCs:
The hardest part is placing the 2x D-sub connectors to scsi on the lemo break out boxes connected to the 110Bs. The breakout boxes can be seen at the very top of the picture Kiwamu took here. These will require a minor modification to the back panel to allow the scsi cable to get out. There are two of these boxes in the new 1X5 rack. These would be connected by scsi to the ADC adapters in the back of the sus IO chassis in 1X4. The connectors are currently behind the new 1X5 rack (along with some spare ADCs/DACs/BOs.
There are 3 cables going from 40 IDC to 37 D-sub (the last 3 wires are not used and do not need to be connected, i.e. 38-40). These plug into the blue and gold ADC adapter box, the top one shown here. There is one spare connection which will remain unused for the moment. The 40 IPC ends plug into the Optical Lever PD boxes in the upper right of the new 1X4 rack (as seen in the top picture here - the boards on the right). At the back of the blue and gold adapter box is a scsi adapter which goes to the back of the IO chassis and plugs into an ADC.
In the back of the IO chassis is a 4th ADC which can be left unconnected at this point. It will eventually be plugged into the BNC breakout box for PEM signals over in the new 1X7 rack, but is unneeded for a sus test.
DACs:
There are 5 cables going from 3 SOS dewhite/anti-image boards and 2 LSC anti-image boards into 3 blue and gold DAC adapter boxes. Currently they plug into the Pentek DACs at the bottom of the new 1X4 rack. Ideally we should be able to simply unplug these from the Pentek DACs and plug them directly into the blue and gold adapter boxes. However at the time we checked, it was unclear if they would reach. So its possible new cables may need to be made (or 40 pin IDC extenders made). These boxes are then connected to the back of the IO chassis by SCSI cables. One of the DAC outputs will be left unconnected for now.
Binary Output:
The Binary output adapter boxes are plugged into the IO chassis BO cards via D-sub 37 cables. Note one has to go past the ADC/DAC adapter board in the back of IO chassis and plug directly into the Binary Output cards in the middle of the chassis. The 50 pin IDC cables should be unplugged from XY220s and plugged into the BO adapter boxes. It is unclear if these will reach.
Timing:
We have a short fiber cable (sitting on the top shelf of the new 1X3 rack) which we can plug into the master timing distribution (blue box located in the new 1X6 rack) and into the front of the SUS IO chassis. It doesn't quite make it going through all the holes at the top of the racks and through the cabling trays, so I generally only plug it in for actual tests.
The IO chassis is already plugged into the c1sus chassis with an Infiniband cable.
So in Summary to plug everything in for a SUS test requires:
- 6x SCSI cables (3 ADC, 3 DAC) (several near bottom of new 1X3 rack)
- 4x 37 D-sub to 37 D-sub connector (end connectors can be found behind new 1X5/1X6 area with the IO chassis stuff - Need to be made) (4 BO)
- 3x 40 IDC to 37 D-sub connectors (end connectors can be found behind new 1X5/1X6 area - Need to be made)(ADC)
- 5x 64 pin ribbon to 40 IDC cable (already exist, unclear if they will reach) (DAC)
- 8x 50 pin IDC ribbon (already exist, unclear if they will reach) (BO)
- 1x Double fiber from timing master to timing card
- 1x Infiniband cable (already plugged in)
Tomorrow, I will finish up a channel numbering plan I started with Kiwamu on Thursday and place it in the wiki and elog. This is for knowing which ADC/DAC/BO channel numbers correspond to which signals. Which ADCs/DACs/BOs the cables plug into matter for the actual control model, otherwise you'll be sending signals to the wrong destinations.
WARNING: The channel numbers on the front Binary Output blue and gold adapter boxes are labeled incorrectly. Channels 1-16 are really in the middle, and 17-32 are on the left when looking at the front of the box. The "To Binary IO Module" is correct. |
3341
|
Sat Jul 31 14:59:33 2010 |
kiwamu | Update | VAC | Re: How to stop and start slow pumpdown | I stopped the pumping at 14:50 pm because I was going back home. I did the same procedure as Koji wrote down (see here).
The P1 pressure reached 32 Torr.
Koji will take over the pumping shift tonight. |
3340
|
Sat Jul 31 10:12:05 2010 |
kiwamu | Update | VAC | Re: How to stop and start slow pumpdown | I resumed the pumping down. It started from 9:55 am. |
3339
|
Sat Jul 31 04:03:11 2010 |
Gopal | Summary | Optic Stacks | Complete Displacement Translational Transfer Function Matrix | Over the past 36 hours, I've run full-fledged FDAs on KALLO.
The transfer functions for translational drives and responses are neatly described by the attached transfer function "matrix."

Next steps:
1) Extend the 3x3 analysis to include tilts and rotations in a 6x6 analysis.
2) Figure out how to do the same types of tests for phase instead of displacement. |
3338
|
Fri Jul 30 22:07:42 2010 |
Koji | HowTo | VAC | How to stop and start slow pumpdown | Kiwamu and Koji
The rough pumping was stopped at 230mmTorr. We are going to resume the pumping tomorrow morning.
- RV1 was closed with torque wheel.
- V3 was closed from MEDM
- RP3 wss stopped from MEDM
- The quick coupling to RP3 was disconnected.
Quote: |
How to stop: 1, close RV1 by torque wheel 2, close V3 from MEDM screen 3, turn off RP3 roughing from MEDM screen 4, disconnect metal hose to oily pump
after butterfly valve. This KF-45 O-ring seal should be kept clean 5, place/close 45 mm cover blanks at the end of the hose and and on the 5" nipple.
How to start: 1, remove blanks from hose and nipple 2, reconnect roughing pump hose to RV1 nipple 3, turn on PR3 4, open V3 5, open RV1 by wrench to ~3/8
6, fine tune RV1 opening to 1 Torr/min
ESSENTIAL: one operator has to be present when oilly roughing pump is connected to the vac. envelope
|
|
3337
|
Fri Jul 30 19:20:15 2010 |
Alastair | Configuration | General | Jacket |
Quote: |
Jenne and I need our jackets. We removed them from the Rb clock.
Thanks for making them warm. Probably a Scotland sweater would fit.
|
Thanks for the loan guys. I do have a lot of warm weather clothing at home that is not so necessary in the California climate. I will find some suitable attire for the Rb clocks there. |
3336
|
Fri Jul 30 17:54:55 2010 |
steve | HowTo | VAC | How to stop and start slow pumpdown |
Quote: |
Bob and Steve closed BS chamber with the help of the manual Genie lift and the pump down started. The PSL shutter was closed and manual block was placed in the beam path. High voltage power supplies were checked to be off.
Pumping speed ~ 1 Torr/min was achieved at 1/8 of a turn opened roughing valve RV1
|
We are at 370 Torr at 9 hrs of pumping. RV1 is opened to ~ 3/8 turn orifice. We are using one roughing pump RP3 and butterfly valve is in place.
How to stop: 1, close RV1 by torque wheel 2, close V3 from MEDM screen 3, turn off RP3 roughing from MEDM screen 4, disconnect metal hose to oily pump
after butterfly valve. This KF-45 O-ring seal should be kept clean 5, place/close 45 mm cover blanks at the end of the hose and and on the 5" nipple.
How to start: 1, remove blanks from hose and nipple 2, reconnect roughing pump hose to RV1 nipple 3, turn on PR3 4, open V3 5, open RV1 by wrench to ~3/8
6, fine tune RV1 opening to 1 Torr/min
ESSENTIAL: one operator has to be present when oilly roughing pump is connected to the vac. envelope |
3335
|
Fri Jul 30 17:52:12 2010 |
Koji | Configuration | General | Jacket | Jenne and I need our jackets. We removed them from the Rb clock.
Thanks for making them warm. Probably a Scotland sweater would fit. |
3334
|
Fri Jul 30 17:24:22 2010 |
Jenne | Update | PEM | Guralps back in the 40m | I brought the GUR2 seismometer back from Bridge so I can get some more MC/Seismic data during the next week while we're pumped down, before we start doing things to the PSL table. Both of the Gur Seismometers are connected back up to the breakout box as of ~3:27pm today. Alastair still has the handheld controller thing (which I use for mass centering, on occasion), since he'll want the seismometer back in a week or two when I'm done with it.
Something is wrong with both X channels of the Guralps. Alastair claimed that he and Frank didn't do anything bad when they opened up the breakout box, but I am suspicious.
While I'm at it, a reminder that Jan and his SURF student Greg still have the Ranger, disassembled over in Bridge. They made a note in their elog, but not in the 40m elog when they took it back again.
Whenever you're done with the Mode Cleaner for the next week, please make sure it is locked, nicely aligned and happy before you leave. Also please make a note of what you're doing and when, so that I know what is good data and what is data with unusual conditions.
In other, semi-bad news (but already recovered from), when I was finishing putting the Guralp Breakout Box back in the rack, I bumped the power strip that is on the top back side of the rack, near the corner that the door opens on (not the corner the door hinges on). I turned the power strip back on, and I think everything that is connected to it came back okay. Anyhow, my bad. Sorry. |
3333
|
Fri Jul 30 14:53:09 2010 |
steve | Update | General | PSL table leg upgrade |
Quote: |
July
29 Thu BS chamber work: Move cable towers / green steering mirrors / (2 TTs with TT charactrization) / Put the heavy door by 5PM.
30 Fri Pumping down
31 Sat WFS work by Nancy
Aug
1 Sun - 5 Thu WFS work by Nancy
5 Thu PSL Table prep
6 Fri PSL Table prep / Likely to shut down the PSL
9 Mon PSL Table prep / shutting down of the PSL (optional)
10 Tue PSL box Frame lifting
12 Thu PSL table tapping
16 Mon - 17 Tue concrete pouring preparation
19 Thu - 23 Fri Tripod placement
24 Tue - 26 Thu concrete pouring
|
The PSL table height will be raised to the level of the AP table. This will be done using 6 TMC's Rigid Damped Tripod legs. Their planed positions shown below.
The legs will be grouted to the floor with concrete.
http://www.techmfg.com/products/floorplatforms/rigiddampedtripod.htm |
Attachment 1: psllegs.PDF
|
|
3332
|
Fri Jul 30 12:46:38 2010 |
Koji | Configuration | 40m Upgrading | BS chamber status | |
Attachment 1: BS_layout_Jul_29_2010.png
|
|
3331
|
Fri Jul 30 09:55:34 2010 |
steve | Configuration | VAC | slow pumpdown has started | Bob and Steve closed BS chamber with the help of the manual Genie lift and the pump down started. The PSL shutter was closed and manual block was placed in the beam path. High voltage power supplies were checked to be off.
Pumping speed ~ 1 Torr/min was achieved at 1/8 of a turn opened roughing valve RV1 |
Attachment 1: slowpd5hr.jpg
|
|
3330
|
Fri Jul 30 09:51:58 2010 |
kiwamu | Update | Green Locking | Re: waist positon of Gaussian beam in PPKTP crystals | Okay, I guess I understand what you want to know. I did the following steps.
1. calculated the conversion efficiency as a function of the waist size. I found w~50um was the optimum waist.
Note: the confocal relation Zr = pi * wo^2/(lamba/n) = L/2 gives us almost the same optimum waist.
elog #2735

2. Did mode matching to get w=50um
elog #2959
elog #3188


3. calculated the offset
elog #2850
4. Moved the ovens
Quote: |
I thought we cared about satisfying the confocal focusing parameter, that is to say we want to set Zr = 2L_crystal. If Zr changes inside the crystal, this is the number we care about..isn't it NOT the waist size, but the rayleigh range we care about? I am not entirely sure what youre response is saying you did...
- Calculate Zr = pi * wo^2/(lamba/n)
- Do mode matching to get this wo in free space
- Calculate the offset you need to move the oven by using n
- Move hte ovens
OR
- Calculate Zr = pi*wo^2/(lamba)
- Do mode matching to get this in free space
- Calculate the offset you need to move your ovens using n
- Move your ovens
I guess the waist size would also let me know - are you using 69 um or 53 um waist size?
|
|
3329
|
Fri Jul 30 02:54:04 2010 |
Koji | Update | 40m Upgrading | More optics installed on the BS table | July 29 Thu [Steve, Alberto, Kiwamu, Koji]
We placed some optics in the BS chamber.
The chambers are ready to be pumped down on Friday once the heavy door is placed.
- Clean room work
- Engraved two Y2 mirrors and PBS@532nm
- Engraved three DLC mounts
- Each of the mounts needs a 3.5 inch post. We found there is no stock of the post in the lab! Also the clamps!
- Took the posts from the temporarily removed optics although we need to return those optics into the table during the next vent.
- We should count the # of the mounts and count the needed posts. Posts and clamps can be either a DLC thick post or New Focus pedestal.
- In the chamber
- The terminal holder was moved as Alberto described
- The green steering optics were placed as Alberto described
- Note: the PBS is flipped in the mount (reflection side is back side)
- The table leveling
- Releasing EQ stops / Check OSEMs / Adjust OSEMs (BS OSEMs are untouched)
- After closing the chamber
Quote: |
[Koji, Steve, Kiwamu, Alberto]
- This afternoon we installed a few new optics on the BS table: GR_PBS, GRY_SM2, GRY_SM1.
- We pulled up the cables so that we had more freedom to move one of the cable towers farther South.
- Then we re-leveled the table. PRM OSEMs were adjusted to be nominal insertions.
- Koji released the earthquake stops on BS but the readout of the OSEMs was apparently frozen on the MEDM screens.
Initially we thought it was a software problem. a nuclear reboot didn't solve it. We spent the following three hours investigating the cause.
Eventually it turned out that the earthquake stops on BS weren't actually fully released.
We opened the tank and accessed to BS. Releasing the earthquake stops in full solved the issue. The OSEMs readout went back to normal.
|
|
3328
|
Fri Jul 30 00:02:15 2010 |
Dmass | Update | Green Locking | waist positon of Gaussian beam in PPKTP crystals |
Quote: |
- As you said, I just calculated the waist position in the crystal because the speed of light changes in a medium and eventually the waist position also changes.
- Yes, I did. Once you get a beam with the right waist size, you just put your crystal at the waist position with the offset.
In fact you don't have to think about the rayleigh range inside of the crystal because what we care is the waist size and it doesn't change.
Quote: |
If I understand your elog, you are just calculating the the offset in position space that you get by having a refractive index.
Did you end up changing the mode matching so that the rayleigh range (which changes with refractive index) was confocally focused inside the crystal (e.g. Zr = 15 mm?
|
|
I thought we cared about satisfying the confocal focusing parameter, that is to say we want to set Zr = 2L_crystal. If Zr changes inside the crystal, this is the number we care about..isn't it NOT the waist size, but the rayleigh range we care about? I am not entirely sure what youre response is saying you did...
- Calculate Zr = pi * wo^2/(lamba/n)
- Do mode matching to get this wo in free space
- Calculate the offset you need to move the oven by using n
- Move hte ovens
OR
- Calculate Zr = pi*wo^2/(lamba)
- Do mode matching to get this in free space
- Calculate the offset you need to move your ovens using n
- Move your ovens
I guess the waist size would also let me know - are you using 69 um or 53 um waist size? |
3327
|
Thu Jul 29 22:58:25 2010 |
kiwamu | Update | Green Locking | waist positon of Gaussian beam in PPKTP crystals | - As you said, I just calculated the waist position in the crystal because the speed of light changes in a medium and eventually the waist position also changes.
- Yes, I did. Once you get a beam with the right waist size, you just put your crystal at the waist position with the offset.
In fact you don't have to think about the rayleigh range inside of the crystal because what we care is the waist size and it doesn't change.
Quote: |
If I understand your elog, you are just calculating the the offset in position space that you get by having a refractive index.
Did you end up changing the mode matching so that the rayleigh range (which changes with refractive index) was confocally focused inside the crystal (e.g. Zr = 15 mm?
|
|
3326
|
Thu Jul 29 22:08:32 2010 |
Alberto | Update | SUS | More optics installed on the BS table | [Koji, Steve, Kiwamu, Alberto]
- This afternoon we installed a few new optics on the BS table: GR_PBS, GRY_SM2, GRY_SM1.
- We pulled up the cables so that we had more freedom to move one of the cable towers farther South.
- Then we re-leveled the table. PRM OSEMs were adjusted to be nominal insertions.
- Koji released the earthquake stops on BS but the readout of the OSEMs was apparently frozen on the MEDM screens.
Initially we thought it was a software problem. a nuclear reboot didn't solve it. We spent the following three hours investigating the cause.
Eventually it turned out that the earthquake stops on BS weren't actually fully released.
We opened the tank and accessed to BS. Releasing the earthquake stops in full solved the issue. The OSEMs readout went back to normal. |
3325
|
Thu Jul 29 21:13:39 2010 |
Dmass | Update | Green Locking | waist positon of Gaussian beam in PPKTP crystals |
Quote: |
The mode profile of Gaussian beams in our PPKTP crystals was calculated.
I confirmed that the Rayleigh range of the incoming beam (1064 nm) and that of the outgoing beam (532 nm) is the same.
And it turned out that the waist postion for the incoming beam and the outgoing beam should be different by 13.4 mm toward the direction of propagation.
These facts will help us making optical layouts precisely for our green locking.
(detail)
The result is shown in the attached figure, which is essentially the same as the previous one (see the entry).
The horizontal axis is the length of the propagation direction, the vertical axis is the waist size of Gaussian beams.
Here I put x=0 as the entering surface of the crystal, and x=30 mm as the other surface.
The red and green solid curve represent the incoming beam and the outgoing beam respectively. They are supposed to propagate in free space.
And the dashed curve represents the beams inside the crystal.
A trick in this calculation is that: we can assume that the waist size of 532 nm is equal to that of 1064 nm divided by sqrt(2) .
If you want to know about this treatment in detail, you can find some descriptions in this paper;
"Third-harmonic generation by use of focused Gaussian beams in an optical super lattice" J.Opt.Soc.Am.B 20,360 (2003)"
|
If I understand your elog, you are just calculating the the offset in position space that you get by having a refractive index.
Did you end up changing the mode matching so that the rayleigh range (which changes with refractive index) was confocally focused inside the crystal (e.g. Zr = 15 mm?
|
3324
|
Thu Jul 29 20:43:32 2010 |
Gopal | Summary | Optic Stacks | Modeling Tips and Tilts | I have discovered a method of completely characterizing the 6x6 response of all six types (x-,y-, and z- translational/rotational) of oscillatory disturbances at the base of the stack.
- "Tipping" drives are trivial, and simply require a face load in the appropriate direction.
- "Tilting" drives could use a torque, but I am instead implementing multiple edge loads in opposing directions to create the appropriate net curl. This curl will be kept constant across the three axes for sake of comparing the resulting transfer functions.
- "Tipping" responses are once again trivial, and merely require the displacement vector of the top center coordinate to be recorded.
- "Tilting" responses require the normal vector to be recorded and manipulated to produce the angular coordinates (assuming right-handed coordinate system):
- θx = tan-1(x/z)
- θy = tan-1(y/z)
- θz = tan-1(y/x)
The first three concepts have been confirmed through simulations to produce correct transfer functions. The last test seems to be producing some problems, in that the vector normal to the equilibrium position (an obvious and useless piece of information) is sometimes given instead of the vector normal to the position of maximum displacement. This means that, as of now, I have the capability of measure the half of the complete 6x6 matrix of transfer functions in the coming weeks. The first three of eighteen transfer functions are attached below and will be included in my progress report.
   |
3323
|
Thu Jul 29 17:12:48 2010 |
josephb, kiwamu | Update | CDS | Re:Working out ADC/DAC/BO wiring | We have installed 4 BO boards, 3 DAC boards and 1 ADC board for new C1SUS.
They are on the 1Y5 rack.
 |
3322
|
Thu Jul 29 17:11:16 2010 |
Gopal | Update | COMSOL Tips | Including Gravity in COMSOL | [Gopal, Jan]
For the past couple of days, Jan and I have been discussing a major issue in COMSOL involving modeling both oscillatory and non-oscillatory forces simultaneously while using FDA. It turns out that he and I had run into the same problem at different times and with different projects. After discussing with an expert, Jan had decided in the past that this simple task was impossible via direct means.
The issue could still be resolved if there was a way for us to work on the Weak Form of the differential equations describing the system:
- Usually, one must define weight as a body load in the negative-z direction. However, this problematically instantiates a new force in COMSOL, which is automatically driven over the range of frequencies during FDA.
- Instead, we could define gravity as an anti-restoring force, since we assume that the base of the stack is fixed.
- In other words, Fg = (ρ*g/L)*x + (ρ*g/L)*y for a point mass which is constrained on the bottom (for small angles).
- Working in Weak Form then, we'd never have to define an explicit gravity load-- this could just be an extra couple of terms in the differential equation which are related entirely to the x- and y-vectors (well-defined for each mesh point). This would fool COMSOL into never tacking on the oscillatory term during FDA.
According to current documentation however, Weak Form analysis is not yet possible in COMSOL 4.0. Jan suggested moving my work over to ANSYS or waiting for the 4.0 upgrade, but there's probably not enough time left in my SURF for either of these options. I suggested attempting a backwards-compatibility test to COMSOL 3.5; Jan and I will be exploring this option some time next week. |
|